the greater of 2 consecutive integers is 15 more than twice the smaller
Answer by checkley77(12569)

LET X & (X+1) BE THE 2 CONSECUTIVE INTEGERS.
X+1=2X+15
X-2X=15-1
-X=14
X=-14 ANS. FOR THE SMALLER INTEGER.
-14+1=-13 ANS. FOR THE LARGER INTEGER.
PROOF:
-13=2*-14+15
-13=-28+15
-13=-13